OpenStudy (anonymous):

WILL FAN AND MEDAL!!!!!!!! 5/3p + 2/3=5+p/2p

jimthompson5910 (jim_thompson5910):

Is the equation this? \[\Large \frac{5}{3}p + \frac{2}{3} = 5 + \frac{p}{2p}\]

OpenStudy (anonymous):

close, the p is with the 3 in the first one

jimthompson5910 (jim_thompson5910):

so \[\Large \frac{5}{3p} + \frac{2}{3} = 5 + \frac{p}{2p}\] ??

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

Are you able to identify the LCD?

OpenStudy (anonymous):

2?

jimthompson5910 (jim_thompson5910):

2 is somewhat part of it

jimthompson5910 (jim_thompson5910):

it's actually 3*2p = 6p

jimthompson5910 (jim_thompson5910):

what would happen if you multiplied both sides by the LCD 6p?

OpenStudy (anonymous):

35p^2/18p^2+12p/18p=35p+6p^2/12p^2 right??

jimthompson5910 (jim_thompson5910):

Multiply both sides by 6p to get \[\Large \frac{5}{3p} + \frac{2}{3} = 5 + \frac{p}{2p}\] \[\Large 6p \left(\frac{5}{3p} + \frac{2}{3}\right) = 6p \left(5 + \frac{p}{2p}\right)\] \[\Large 6p\left(\frac{5}{3p}\right) + 6p\left(\frac{2}{3}\right) = 6p \left(5\right) + 6p \left(\frac{p}{2p}\right)\] \[\Large \frac{30p}{3p} + \frac{12p}{3} = 30p + \frac{6p^2}{2p}\] \[\Large 10 + 4p = 30p + 3p\] Notice how the fractions are now gone.

jimthompson5910 (jim_thompson5910):

Do you see how I'm getting that?

OpenStudy (anonymous):

sort of..

jimthompson5910 (jim_thompson5910):

Which step is confusing you?

OpenStudy (anonymous):

nvm, i get it. so now what?